# An Adaptive Placement and Parallelism Framework for Accelerating RLHF Training

### v2

@@ III-B Key Challenges
To overcome these limitations, it is crucial to explore new distributed model placement strategies for efficient RLHF training, which also could save substantial costs for large-scale training prevalent in this scenario. However, to build an efficient and easy-to-use RLHF training framework, we encounter multiple challenges, including:
@@
Trade-off Between Memory and Communication in Complex System. The design of the Model Placement strategy for the RLHF system presents complexity from two perspectives. Firstly, it requires considering the intricate dependencies among each model in RLHF training. Failure to do so can result in unnecessary data interaction and device idleness aforementioned. Secondly, it needs careful trade-offs between memory and communication costs for the collective set of four models in either training or inference mode, as well as for each individual model. This necessitates considering different parallelism strategies or optimization techniques based on the scale of model size and devices.
@@ IV Our Framework
Therefore, we introduce the Adaptive Placement and Parallelism (APP) Framework for accelerating the RLHF Training. Firstly, besides the Flattening strategy, we propose two model placement strategies aimed at enhancing the efficiency of RLHF training. These strategies are carefully designed to optimize communication efficiency and utilize memory effectively, catering to different training scenarios that involve different model scales and hardware heterogeneity. Secondly, we present the design of an APP Execution Engine abstraction, which provides modeling experts with a simplified programming API by decoupling RLHF modeling from distributed computing.
@@ IV-A1 Interleaving Strategy
The main idea of the Interleaving strategy is to partially parallelize the pipeline by allocating the independent models on exclusive group of GPU devices, as well as shrinking the number of computing devices required for each model. We observe that either Reward&Ref Inference (and Actor&Critic training models) could operate independently, thereby it is possible to eliminate part of the serial execution. For example, the Ref model and Reward model both take the generated responses from the Actor model as input and output the intermediate results into the experience buffer without synchronization. Also, training Actor and Critic nodes have no dependency since they do not need synchronization during the training phase. This design helps the parallelization of independent models in RLHF training. This lowers the memory redundancy and shrinks the scale of participating devices used by independent models rather than using all devices in the Flattening strategy.
@@
We compare the distributed architecture and computation procedure between the Flattening strategy and our Interleaving strategy. We use a small number of devices as illustrative examples, while our strategy could be easily extended to large-scale computing devices. In terms of Model Placement, the Flattening strategy treats the four models as a unit and deploys them on all devices, as depicted in Figure 2(a). By contrast, our Interleaving strategy distinctly assigns the Ref model and Reward model to two distinct device groups (GPU 0-3 and 4-7)222Our interleaving strategy could further place the Actor and Critic model on exclusive devices. We only illustrate the interleaving between Reward and Ref for simplicity., as illustrated in Figure 2(b). It reduces the number of devices by half used by either the Ref model or the Reward model. It’s important to note that responses generated from the Actor are located on GPUs where the Ref model is not placed since the Actor model is deployed on all GPU devices, as shown in the timeline in Figure 3(b). Therefore, it is necessary to perform two extra communication stages: an AllGather primitive to gather the generated results from all devices before forwarding, and then perform a Scatter primitive to distribute the results back to all devices. However, these two additional communication stages can be considered negligible by overlapping with computation. In other words, computation and communication occupy different cuda streams, allowing computation to begin upon receiving the first batch while the remaining batches can be communicated synchronously with the ongoing computation. We conclude that the Interleaving strategy shortens the time-cost generation duration from two aspects.
@@
Firstly, in terms of memory consumption, our Interleaving strategy reduces the number of participating devices for either the Ref or Reward model, effectively minimizing memory redundancy when applying ZeRO Parallelism. This allows our method to leverage the saved memory to increase the total batch sizes, resulting in higher overall throughput. For instance, when employing the Flattening strategy as displayed in Figure 3(a) under ZeRO-0, all devices hold replicas of the model states, e.g. of Reward Model, while fewer devices need to hold the replicas under our Interleaving strategy as depicted in Figure 3(b).
@@
Secondly, from a communication efficiency perspective, having fewer participating devices leads to reduced communication overhead in both the generation and training stages. It’s important to note that while increased levels of ZeRO parallelism reduce memory redundancy in model states, it comes at the cost of increased communication as discussed in SectionII-B. In this case, the Interleaving strategy can further leverage high-speed intra-node bandwidth and reduce inter-node communication by colocating the Reward model or Ref model on a single machine with usually 8 GPU devices, particularly in small to medium training scenarios, e.g. 7B or 13B models.
IV-A2 Separation Strategy
@@
To shorten the generation duration, our Separation strategy involves separating the training and the inference runtime of the RLHF pipeline. Specifically, we add extra shadow actor and critic models for inference only without a mixture usage of runtime. These shadow models are dedicated to inference and their parameters are synchronized with the corresponding training models periodically. This division not only allows us to run the training and generation models simultaneously but also leverages optimization techniques tailored for training or inference. It also could easily utilize the extra heterogeneous devices in clusters, since the high-end homogeneous GPU devices are scarce. Consequently, this approach significantly accelerates the generation stage and re-balance the workloads between these stages in the RLHF.
@@
This strategy significantly speeds up the entire RLHF training, though it introduces certain memory redundancy due to additional models and communication overhead. Figure 2(c) illustrates the placement of the training Actor and Critic models and inference models, including the Shadow Actor and Critic models, on separate devices. Optionally, we could further place the training Actor and Critic models on exclusive devices in an interleaving manner to further decrease the memory redundancy and communication costs. However, there is a need for communication for intermediate results and parameter synchronization between the training Actor/Critic models and the prediction models. Firstly, as shown in Figure 4(a), extra point-to-point communications are required before or after the Forward stage of each prediction model. Additionally, the training Actor and Critic models need to pull the intermediate results from the inference devices. These extra communication stages introduce some idle time per device, known as bubble overhead. However, these bubbles can be minimized by overlapping computation and communication stages through the use of pipeline parallelism as discussed in Section II-A. As illustrated in Figure 4(a), the bubbles are nearly hidden through micro-batching in a pipeline parallelism manner, especially with a higher number of generation times. Secondly, it is necessary to synchronize the trained model parameters between the training Actor/Critic models and shadow models after each round of PPO training. The synchronization overhead is inevitable but acceptable333It is critical to clarify that this process is not asynchronous training, but still synchronous training without compromising model accuracy. It thanks to the off-policy property of RLHF that the training phase involves sampling from a buffer rather than tightly following the forward stage..
@@
Although this approach introduces some complexity and overhead, it provides significant benefits from three different perspectives: 1) Our Separation strategy decouples training and inference runtime of the pipeline, enabling targeted optimization techniques for each stage, as discussed in Section III-A. For example, the time-consuming Actor Generation can be accelerated by utilizing specialized parallelism strategies that are well-suited for inference workloads, such as intra-node tensor parallelism by reducing low-bandwidth inter-node communication as discussed in II. 2) The strategy also facilitates the integration of evolving inference optimizations like quantization[34], which are less compatible with model training. There will be considerable runtime swapping overheads when using a mixture of training and inference runtime. 3) The strategy allows scaling out additional heterogeneous devices tailored for training or inference workloads, especially for the time-consuming generation stage. This flexibility in either quantity or type of devices offers significant adaptability to diverse training scenarios.
IV-B APP Execution Engine
@@
The APP Execution Engine is designed to provide a user-friendly interface for modeling experts to easily build their own RLHF training pipeline while hiding the messy details of distributed computing and intermediate data interaction as illustrated in Figure 5. The APP Execution Engine encapsulates each model involved in the RLHF pipeline considering both AC-Share and AC-NonShare structures. Our modular design allows for the encapsulation of each model and corresponding distributed techniques. These techniques include the model placement strategy, singular model parallelism (such as ZeRO in DeepSpeed and Tensor Parallelism in Megatron), as well as various optimization techniques in other libraries. We further the abstraction in detail:
@@
Model Placement Abstraction. While our placement strategies offer general guidelines on how to distribute these models, there remains a gap in determining the optimal number of devices for deploying a target model, as this relies on factors like model sizes and available hardware in different training scenarios. To address this, we introduce the concept of Model Placement Ratio, which is a value ranging from 0 to 1 that represents the proportion of devices allocated for a specific model. This provides enhanced adaptability. For instance, the Flattening strategy for a Reward model has a ratio of 1, indicating that the model is deployed on all devices. On the other hand, the Interleaving strategy for the Reward model, as depicted in Figure 3, has a ratio of 0.5, signifying that the model is deployed on half of the devices. Take the Reward Model in the model placement ratio representation shown in Figure 5, e.g. [1, 1, 0.5, 0.5], as an example. The engine constructs a communication subgroup on the deployed devices of the Reward model for synchronization. It also belongs to a parent communication group that is connected with all models’ subgroups for data interaction. Then our Execution Engine generates the model placement mapping, which places the model to physical devices properly according to the model placement ratio, considering the device capacity and topology, such as intra- and inter-node. This mapping involves detailed ranking assignments for each model.
@@
Hiding Intermediate Data Interaction. Data interaction caused by the placement strategy is also encapsulated within the Execution Engine. The Interleaving and Separation strategies introduce communication among different communication subgroups during the different stages as discussed in Section IV-A, and the APP engine handles these data interaction logics. Each training or forward operation of the models is treated as an individual Operator. The rank assigned to each model performs an AllGather operation to collect the upstream data before executing the Operator. Once the Operator is executed, the resulting data is scattered back downstream.
@@
Overall, this abstraction enables a general interface and backend for creating a wide range of RLHF algorithms for research or industrial applications. Users can focus on the essence of RLHF modeling without being burdened by the intricacies of the underlying details, as shown in Figure 5.
@@ V Experiments
We conduct comparative experiments on both AC-NonShare and AC-Share model structures against the Flattening strategy used by DeepSpeed-Chat and trlX respectively. We evaluate the model throughput while using the maximum batch size available for each framework before Out-of-Memory (OOM) for fairness. These experiments encompass variations in model sizes, device scales, and device types, allowing us to assess the effectiveness and adaptability of our APP framework across different training scenarios.
V-A Experiment Settings
LLM Backbones. We employ Llama[26] of different parameter sizes as our LLM backbone, including 7B, 13B, 33B, and 65B. For demonstration purposes, we keep consistent model size for each model involved in RLHF pipeline, but our APP framework could adapt to training scenarios where each RLHF model may have varying model sizes and other LLM models, such as Opt[35].
Hyperparameters. We have established the following parameters for our research. Both the number of PPO training epochs and batches per rollout are set to 1. The input sequence length is limited to 256. Both minimum and maximum generation lengths are 256. The batch size for each experiment is set to the maximum value before Out of Memory. For other model convergence-related parameters, we keep consistent for fairness.
@@ V-A Experiment Settings
Baseline. We conduct a comparison between the APP framework and the current state-of-the-art methods in RLHF training: DeepSpeed-Chat[33] with Hybrid Engine for the AC-NonShare case and trlX[3] for the AC-Share case.
Evaluations. We evaluate the effectiveness of different training frameworks by comparing their sample throughput during training. Sample throughput refers to the rate at which samples are processed end-to-end, typically measured in samples per second(# samples/sec). Furthermore, we verify that the model convergence, measured in returns, remains unaffected by our placement strategies, as evidenced in Figure 9, when compared to the DeepSpeed-Chat implementation.
@@ V-B Experiments and Analysis
We compare the throughput of our APP framework against DeepSpeed-Chat and trlX respectively. Specifically, in the training of the 7B model, we utilize ZeRO-2 for training models and DDP for inference models. For the training of more large models, e.g., 13B, 33B, and 65B models, we have to employ ZeRO-3 for both stages. Furthermore, our Separation strategy incorporates tensor parallelism.
V-B1 Comparing with DeepSpeed-Chat

### v3

@@ III-C Key Challenges
To overcome these limitations, it is crucial to explore new distributed model placement strategies for efficient RLHF training, which could also save substantial costs for large-scale training prevalent in this field. However, to build an efficient and easy-to-use RLHF training framework, we encounter multiple challenges, including:
@@
Trade-off Between Memory and Communication in Complex System. The design of the Model Placement strategy for the RLHF system presents complexity from two perspectives. Firstly, it requires considering the intricate dependencies among each model in RLHF training. Failure to do so can result in unnecessary data interaction and device idleness as aforementioned in trlX. Secondly, it needs careful trade-offs between memory and communication costs for the collective set of four models in either training or inference mode, as well as for each individual model. This necessitates considering different parallelism strategies or optimization techniques based on the scale of model size and hardware heterogeneity.
@@ IV Our Framework
Therefore, we introduce our FlexRLHF Framework for accelerating the RLHF Training. Firstly, besides the Co-located strategy, we propose two model placement strategies aimed at improving the efficiency of RLHF training. These strategies are carefully designed to optimize communication efficiency and utilize memory effectively, catering to different training scenarios that involve different model scales and hardware heterogeneity. Secondly, we present the design of a FlexRLHF Execution Engine abstraction, which provides modeling experts with a simplified programming API by decoupling RLHF modeling from distributed computing. Lastly, we offer a guideline on how to configure the distributed strategies for efficient RLHF training in practice.
Input: devices; Actor, Critic, Reward, and Ref models
@@ IV-A1 Interleaving Strategy
The main idea of the Interleaving strategy is to partially parallelize the pipeline by allocating the independent models to exclusive group of GPU devices, as well as shrinking the number of computing devices required for each model. We observe that either Reward&Ref Forward (and Actor&Critic training models) could operate independently, thereby it is possible to eliminate part of the serial execution. For example, the Ref model and Reward model both take the generated responses from the Actor model as input and output the intermediate results into the experience buffer without synchronization. Also, training Actor and Critic nodes have no dependency since they do not need synchronization during the training phase. This design helps the parallelization of independent models in RLHF training. This lowers the memory redundancy and shrinks the scale of participating devices used by independent models rather than using all devices in the Co-located strategy.
@@
We compare the distributed architecture between the Co-located strategy and our Interleaving strategy using a small number of devices as illustrative examples shown in Fig. 2. Following Algorithm 1, our strategy could be easily extended to large-scale computing devices. In terms of Model Placement, the Co-located strategy treats the four models as a unit and deploys them on all devices, as depicted in Fig. 2(a) and Fig. 4. By contrast, our Interleaving strategy distinctly assigns the Ref model and Reward model to two distinct device groups (GPU 0-3 and 4-7)222Our interleaving strategy could further place the Actor and Critic model on exclusive devices. We only illustrate the interleaving between Reward and Ref for simplicity., as illustrated in Fig. 2(b). It reduces the number of devices by half used by either the Ref model or the Reward model. It’s important to note that responses generated from the Actor are located on GPUs where the Ref model is not placed, since the Actor model is deployed on all GPU devices, as shown in the timeline in Fig. 4(b). Therefore, it is necessary to perform two extra communication stages: an AllGather primitive to gather the generated results from all devices before the forward stage, and then an AlltoAll primitive to distribute the results over all devices. However, these two additional communication stages can be considered negligible by overlapping with computation. In other words, computation and communication occupy different cuda streams, allowing computation to begin upon receiving the first batch while the remaining batches can be communicated synchronously with the ongoing computation. We conclude that the Interleaving strategy shortens the time-cost generation duration from two aspects:
@@
Firstly, in terms of memory consumption, our Interleaving strategy reduces the number of participating devices for either the Ref or Reward model, effectively minimizing memory redundancy when applying DP-based Parallelism, e.g., ZeRO. This allows our method to leverage the saved memory to increase the total batch sizes, resulting in higher overall throughput. For instance, when employing the Co-located strategy under ZeRO-0 as displayed in Fig. 4(a), all devices hold replicas of the model states, e.g. of Reward Model, while half of devices need to hold the replicas under our Interleaving strategy as depicted in Fig. 4(b).
@@
Secondly, from a communication efficiency perspective, having fewer participating devices leads to reduced communication overhead in both the generation and training stages. Also, it’s important to note that while increased levels of ZeRO parallelism reduce memory redundancy in model states, it comes at the cost of increased communication as discussed in Section II-B. In this case, the Interleaving strategy can further leverage high-speed intra-node bandwidth and reduce inter-node communication by colocating the Reward model or Ref model on a single machine, particularly in small to medium training scenarios, e.g. 7B or 13B models.
Input: training devices, and inference devices; Actor, Critic, Reward, and Ref models.
Running following processes on inference devices concurrently
Running following processes on training devices concurrently
IV-A2 Disaggregated Strategy
@@
The main idea of our Disaggregated strategy is to physically disaggregate the training and the inference runtime of the RLHF pipeline to shorten the generation duration. Specifically, we add extra actor and critic replicas as shadow models for inference only without a mixture usage of runtime as shown in Fig. 2(c). These shadow models are dedicated to inference and their parameters are synchronized with the corresponding training models periodically. This division not only allows us to run the training and generation stages simultaneously but also leverages optimization techniques tailored for training or inference. It also could easily utilize the extra heterogeneous devices in clusters, since the high-end homogeneous GPU devices are scarce. Overall, this approach significantly accelerates the generation stage and re-balance the workloads among these stages in the pipeline.
@@
Following Algorithm 2, we classify these devices into two groups: training or inference only. This strategy places training models (Actor and Critic models) and inference models, including the Shadow Actor and Critic models, on separate devices. Firstly, we create the communication groups for either inference or training models for further data communication. Secondly, as shown in Fig. 4, extra point-to-point communications are required before or after the Forward stage of each prediction model. Additionally, the training Actor and Critic models need to pull the intermediate outputs from the inference devices. These extra communication stages introduce some idle time per device, known as bubble overhead. However, these bubbles can be minimized by overlapping computation and communication stages in a pipeline manner. As illustrated in Fig. 4(a), the bubbles are nearly hidden through micro-batching using the pipeline technique. Secondly, it is necessary to synchronize the trained model parameters between the training Actor/Critic models and shadow models after each round of PPO training. The synchronization overhead is inevitable but acceptable333It is critical to clarify that this process is not asynchronous training, but still synchronous training without compromising model accuracy. It leverages the off-policy property of PPO that the training phase involves sampling from a buffer rather than tightly following the forward stage.. Although this approach introduces some complexity and overhead, it provides significant benefits from three different perspectives:
@@
1) Our Disaggregated strategy decouples training and inference runtime of the pipeline, enabling targeted optimization techniques for each stage, as discussed in Section III-B. For example, the time-consuming Actor Generation can be accelerated by utilizing specialized framework or parallelism strategies that are well-suited for inference workloads and avoid AllGather overhead compared with using a Hybrid Engine. 2) The strategy also facilitates the integration of evolving inference optimizations, like PagedAttention. There will be considerable runtime swapping overheads when using a mixture of training and inference runtime. 3) The strategy allows scaling out additional heterogeneous devices tailored for training or inference workloads, especially for the time-consuming generation stage. This flexibility in either quantity or type of devices offers significant adaptability to diverse training scenarios in practice.
IV-B FlexRLHF Execution Engine
@@
The FlexRLHF Execution Engine is designed to provide a user-friendly interface for modeling experts to easily build their own RLHF training pipeline while hiding the messy details of distributed computing and intermediate data interaction as illustrated in Fig. 5. The FlexRLHF Execution Engine encapsulates each model involved in the RLHF pipeline considering both AC-Share and AC-NonShare structures. Our modular design allows for the encapsulation of each model and corresponding distributed techniques. These techniques include the model placement strategy, singular model parallelism (such as data parallelism in DeepSpeed and model parallelism in Megatron). We further explain the abstraction in detail.
@@
Model Placement Abstraction. To provide a simple user interface to configure all these strategies, we introduce the concept of Model Placement Ratio, which is a value ranging from 0 to 1 that represents the proportion of devices allocated for a specific model. This offers enhanced adaptability. For instance, the Co-located strategy for a Reward model has a ratio of 1, indicating that the model is deployed on all devices. On the other hand, the Interleaving strategy for the Reward model has a ratio of 0.5, signifying that the model is deployed on half of the devices, as depicted in Fig. 4. Take the Reward Model in the model placement ratio representation shown in Fig. 5, e.g. [1, 1, 0.5, 0.5], as an example. The engine constructs a communication subgroup on the deployed devices of the Reward model for synchronization. It also belongs to a parent communication group that is connected with all models’ subgroups for data interaction. Then our Execution Engine generates the model placement mapping, which places the model to physical devices properly according to the model placement ratio, considering the device capacity and topology, such as intra- and inter-node. This mapping involves detailed rank assignments for each model.
@@
Hiding Intermediate Data Interaction. Data interaction caused by the placement strategy is also encapsulated within the Execution Engine. The Interleaving and Disaggregated strategies introduce communication among different communication subgroups during the different stages as discussed in Section IV-A, and the FlexRLHF engine handles these data interaction logics. Each training or forward operation of the models is treated as an individual Operator. The communication group performs an AllGather operation to collect the upstream data before executing the Operator. Once the Operator is executed, the resulting data is scattered back downstream.
@@
Overall, this abstraction enables a general interface and backend for different RLHF algorithms beyond PPO, for research or industrial applications. Users can focus on the essence of RLHF modeling without being burdened by the intricacies of the underlying details.
IV-C A Guideline for Efficient Training
The guideline addresses the following question: “Given devices and the scale of models, how does one determine the optimal configurations of devices or placement ratio for each model?” While our proposed placement strategies improve the efficiency and adaptability of RLHF training, there remains a gap in determining the optimal number of devices for each model under different training scenarios where model sizes and the scale of GPU devices vary. Distributed training experts could easily employ our model placement ratio abstraction for configuring these strategies. Yet, for modeling experts lacking distributed computing expertise, misconfiguration may downgrade the training efficiency as well as incur frequent occurrences of Out-Of-Memory errors. This difficulty originates from the vast hyperparameter space in the distributed RLHF training pipeline. This often involves coordination between these four or six models and their respective distributed strategies. To address these problems, we introduce the following guidelines to assist modeling experts in the configuration process for more efficient distributed training.
Firstly, in scenarios where computing resources are limited, it is recommended to employ the Interleaving strategy and the Co-located strategy. The interleaving strategy is suitable for cases when either the Ref or Reward model can be accommodated within the memory limits of a single node, including multiple GPU devices. In other cases, the Co-located strategy is advised. This is because inter-node communication is more costly for an Interleaving Strategy than the Co-located strategy, and the Disaggregated strategy requires more memory. However, the usage of the Co-located strategy should be infrequent when computing resources are enough since the Disaggregated strategy becomes a more favorable choice in configurations involving more than two nodes. Furthermore, it is advisable to increase the batch size as long as there is spare memory capacity, stopping before 95% memory capacity.
Secondly, employing the Disaggregated strategy proves advantageous when computational resources are enough. 1) To expedite the time-intensive generation phase, it is recommended to allocate additional devices to the Shadow Actor model, provided the minimal memory prerequisites of other models have been satisfied. Guided by our experience, we suggest assigning between 30% to 50% of the total computational devices exclusively to the Actor model for generation purposes. 2) Utilizing multiple replicas of the Actor model in a DP manner is beneficial, particularly when inter-node communication is inevitable, as it helps mitigate high communication costs. For instance, employing a replica of the Actor model on each node is preferable to distributing the model over both in scenarios involving two nodes. 3) Furthermore, to optimize throughput using pipeline techniques, it is advisable to configure the model placement ratio for each model to achieve a uniform distribution of processing time across the generation, forward, and training stages. This reduces the bubble overhead of idle devices in different micro-batches and enables stages to work on distinct micro-batches concurrently.
@@ V Experiments
We conduct comparative experiments on both AC-NonShare and AC-Share model structures against the Co-located strategy used by DeepSpeed-Chat and trlX respectively. These experiments encompass variations in model sizes, device scales, and device types, allowing us to assess the effectiveness and adaptability of our framework across different scenarios.
V-A Experiment Settings.
LLM Backbones. We employ Llama [35] of different parameter sizes as our LLM backbone, including 7B, 13B, 33B, and 65B. For demonstration purposes, we keep consistent model size for each model involved in RLHF pipeline, but our FlexRLHF framework could adapt to training scenarios where each RLHF model may have varying model sizes and other LLM models, such as GPT [4], and Qwen [42].
Hyperparameters. The batch size for each experiment is set to the maximum value before Out of Memory. Both the number of PPO training epochs and the number of batches per rollout are set to 1. The input sequence length is 256 and LoRA dim is 4. Both minimum and maximum generation lengths are 256. For other model convergence-related parameters, we keep consistent for fairness.
@@
Baseline. We conduct a comparison between the FlexRLHF framework and the current state-of-the-art methods in RLHF training: DeepSpeed-Chat [43] with Hybrid Engine for the AC-NonShare case and trlX [12] for the AC-Share case.
Evaluations. We evaluate the effectiveness of different training frameworks by comparing their sample throughput during training. Sample throughput refers to the rate at which samples are processed end-to-end, typically measured in samples per second(# samples/sec). Furthermore, we verify that the model convergence, measured in returns, remains unaffected by our placement strategies, as evidenced in Fig. 9, when compared to the DeepSpeed-Chat implementation.
@@ V-B Experiments and Analysis
We compare the throughput of our framework against DeepSpeed-Chat and trlX respectively. Specifically, in the training of the 7B model, we utilize ZeRO2 for training models and DP for inference models. For the training of more large models, e.g., 13B, 33B, and 65B models, we have to employ ZeRO3 for both stages. Furthermore, our Disaggregated strategy leverages tensor and data parallelism in Megatron.
V-B1 Comparison in AC-NonShare Structure